A sphere of radius 1.00 cm is placed in the path of a parallel beam of light of large aperture. The intensity of the light is 0.50 W/cm2. If the sphere completely absorbs the radiation falling on it, find the force exerted by the light beam on the sphere.

Grade:12

1 Answers

Arun
6 years ago
 

Change in momentum of photon due to absorbtion  =  h /λ 

 

no of photons per unit time   =  p/(hc/λ )      p  = power of beam

 

therefore change in momentum per unit time  =  force  =  h /λ  * p/(hc/λ )      =  p /c 

 

p = intensity * area =  .5 * 4*pie*r2  =   .5 *4*3.14*(1)2  =  6.28 W

 

so force  =  6.28 /3 *108 =  2.09 * 10-8 N . The answer of this question in some books as found by me is incorrect